OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)?

OpenStudy (anonymous):

find the gradient first then sub in any of the point

OpenStudy (anonymous):

4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (anonymous):

First) find slope: y - y / x - x 13 - 1= 12 -2 - 1 = -3 slope = 12/3 = 4/1 = 4 Find Point slope form: y - y = m(x - x) y - 1 = 4(x - 1) y - 1 = 4x - 4 +1 y = 4x - 3 in standard form is: y - 4x = -3

OpenStudy (anonymous):

Hope this helps!